OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).

OpenStudy (anonymous):

First you have to find the slope of that line it is 2 \[y-3=2(x-5)\] \[y-2x+7=0\] it is the equation